# An EFT for anisotropic anti-ferromagnets: gapped Goldstones, pseudo-Goldstones, and phase transitions

### v1

@@ Appendix B Overlap functions from “polology”
To verify our procedure, we re-derive the overlap functions with an alternative method, using the so-called “polology” technique Weinberg (2005), in a way analogous to what recently done in Hui et al. (2024); Creminelli et al. (2024). As usual for such kind of manipulations, we will be using the completeness relation,
$\mathds{1}=\sum_{\alpha=\pm}\int\frac{d^{3}q}{(2\pi)^{3}2\omega_{{q},\alpha}}\ket{\bm{q},\alpha}\bra{\bm{q},\alpha}+\dots\,,$ (40)
@@
where the ellipsis stands for multi-particle states. The latter will not play any role as long as one is interested in tree level $S$-matrix elements, for which one considers the on-shell limit of the external one-particle states. Inserting the above relation in the time-ordered two-point function of the magnon field, one gets,
$\begin{split}\braket{0}{\mathcal{T}\left\{{\theta^{a}}(\bm{x},t){\theta^{b}}(\bm{y},t^{\prime})\right\}}{0}={}&\braket{0}{{\theta^{a}}(\bm{x},t){\theta^{b}}(\bm{y},t^{\prime})}{0}\theta(t-t^{\prime})+\braket{0}{{\theta^{b}}(\bm{y},t^{\prime}){\theta^{a}}(\bm{x},t)}{0}\theta(t^{\prime}-t)\\ ={}&\sum_{\alpha=\pm}\int\frac{d^{3}q}{(2\pi)^{3}2\omega_{q,\alpha}}\big{[}\langle 0|\theta^{a}(\bm{x},t)|\bm{q},\alpha\rangle\langle\bm{q},\alpha|\theta^{b}(\bm{y},t^{\prime})|0\rangle\theta(t-t^{\prime})\\ &+\langle 0|\theta^{b}(\bm{y},t^{\prime})|\bm{q},\alpha\rangle\langle\bm{q},\alpha|\theta^{a}(\bm{x},t)|0\rangle\theta(t^{\prime}-t)\big{]}+\dots\\ ={}&\int\frac{d^{4}q}{(2\pi)^{4}}e^{-i\omega(t-t^{\prime})+i\bm{q}\cdot(\bm{x}-\bm{y})}\sum_{\alpha=\pm}\frac{1}{2\omega_{q,\alpha}}\left[\frac{i\mathcal{Z}^{a}_{q,\alpha}\mathcal{Z}^{b*}_{q,\alpha}}{\omega-\omega_{q,\alpha}+i\varepsilon}-\frac{i\mathcal{Z}^{a*}_{q,\alpha}\mathcal{Z}^{b}_{q,\alpha}}{\omega+\omega_{q,\alpha}-i\varepsilon}\right]+\dots\,,\\ \equiv{}&\int\frac{d^{4}q}{(2\pi)^{4}}e^{-i\omega(t-t^{\prime})+i\bm{q}\cdot(\bm{x}-\bm{y})}\mathcal{K}^{ab}(\omega,q)+\dots\,,\end{split}$ (41)
where $d^{4}q=d\omega d^{3}q$, |$\ket{0}$ is the vacuum in the presence of external magnetic field and anisotropies (see Section II.2), and we used the identity $\theta(t)=\int\frac{d\omega}{2\pi}\frac{i}{\omega+i\varepsilon}e^{-i\omega t}$.
@@
[…]
×$\times\Big{(}\delta^{ab}\big{[}v_{\theta}^{2}q^{2}-\mu^{2}B^{2}-\omega^{2}+{2(\lambda_{z}+\lambda_{z})}\big{]}$
$\quad-\delta^{a2}\,\delta^{b2}{2\lambda_{x}}-2i\mu B\omega\,\epsilon^{ab}\Big{)}\,.$
@@
$\sum_{\alpha=\pm}|\mathcal{Z}_{q,\alpha}^{1}|^{2}\omega_{q,-\alpha}^{2}=v_{\theta}^{2}q^{2}-\mu^{2}B^{2}+{2(\lambda_{z}+\lambda_{z})}\,,$ (43)
@@
$\sum_{\alpha=\pm}|\mathcal{Z}_{q,\alpha}^{2}|^{2}\omega_{q,-\alpha}^{2}=v_{\theta}^{2}q^{2}-\mu^{2}B^{2}+{2\lambda_{z}}\,.$ (44)
Therefore we find that,
${|\mathcal{Z}^{1}_{q,\pm}|}={}$ $\sqrt{\frac{1}{2}\pm\frac{2\mu^{2}B^{2}-\lambda_{x}}{2\sqrt{4\mu^{2}B^{2}(v_{\theta}^{2}q^{2}+\lambda_{x}-2\lambda_{z})+\lambda_{x}^{2}}}}\,,$
${|\mathcal{Z}^{2}_{q,\pm}|}={}$ $\sqrt{\frac{1}{2}\pm\frac{2\mu^{2}B^{2}+\lambda_{x}}{2\sqrt{4\mu^{2}B^{2}(v_{\theta}^{2}q^{2}+\lambda_{x}-2\lambda_{z})+\lambda_{x}^{2}}}}\,,$
in agreement with Eqs. (27) and (28). To determine the relative phase between $\mathcal{Z}^{1}_{q,\alpha}$ and $\mathcal{Z}^{2}_{q,\alpha}$ we write them as,
@@
By taking $a=1$, $b=2$ in Eq. (B), we obtain the following equation for the relative phase, $\Delta\varphi_{\alpha}\equiv\varphi_{\alpha}^{1}-\varphi_{\alpha}^{2}$,
$\begin{split}&\sum_{\alpha}\frac{i|\mathcal{Z}_{q,\alpha}^{1}||\mathcal{Z}_{q,\alpha}^{2}|}{\omega_{q,\alpha}\left(\omega^{2}-\omega_{q,\alpha}^{2}\right)}\left[i\,\omega\sin\Delta\varphi_{\alpha}+\omega_{q,\alpha}\cos\Delta\varphi_{\alpha}\right]\\ &\hskip 85.35826pt=-\frac{2\mu B\omega}{\det\left[\mathcal{M}(\omega,\bm{q})\right]}\,.\end{split}$ (46)

### v2

@@ Appendix C Overlap functions from “polology”
To verify our procedure, we rederive the overlap functions with an alternative method, using the so-called “polology” technique [39], in a way analogous to what has been recently done in Ref. [31, 34]. As usual for such kind of manipulations, we will be using the completeness relation,
$|+,\displaystyle\mathds{1}=\sum_{\alpha=\pm}\int\frac{d^{3}q}{(2\pi)^{3}2\omega_{{q},\alpha}}\ket{\bm{q},\alpha}\bra{\bm{q},\alpha}$$\dots\,$, (42)
@@
where the ellipsis stands for multi-particle states. The latter will not play any role as long as one is interested in tree level $S$-matrix elements, for which one considers the on-shell limit of the external one-particle states. Inserting the above relation in the time-ordered two-point function (or, alternatively, the correlator) of the magnon field, one gets
[…]
@@
$={}$ $\frac{-i}{\det\left[\mathcal{M}(\omega,{q})\right]}$ (44)
$\times\Big{(}\delta^{ab}\big{[}v_{\theta}^{2}q^{2}-\mu^{2}{\rm H}^{2}-\omega^{2}+{2(\lambda_{z}+\lambda_{z})}\big{]}$
$\quad-\delta^{a2}\,\delta^{b2}{2\lambda_{x}}-2i\mu{\rm H}\omega\,\epsilon^{ab}\Big{)}\,.$
@@
$\sum_{\alpha=\pm}|\mathcal{Z}_{q,\alpha}^{1}|^{2}\omega_{q,-\alpha}^{2}=v_{\theta}^{2}q^{2}-\mu^{2}{\rm H}^{2}+{2(\lambda_{z}+\lambda_{z})}\,$, (45)
@@
$\sum_{\alpha=\pm}|\mathcal{Z}_{q,\alpha}^{2}|^{2}\omega_{q,-\alpha}^{2}=v_{\theta}^{2}q^{2}-\mu^{2}{\rm H}^{2}+{2\lambda_{z}}\,.$ (46)
Therefore we find that
${|\mathcal{Z}^{1}_{q,\pm}|}={}$ $\sqrt{\frac{1}{2}\pm\frac{2\mu^{2}{\rm H}^{2}-\lambda_{x}}{2\sqrt{4\mu^{2}{\rm H}^{2}(v_{\theta}^{2}q^{2}+\lambda_{x}-2\lambda_{z})+\lambda_{x}^{2}}}}\,$,
${|\mathcal{Z}^{2}_{q,\pm}|}={}$ $\sqrt{\frac{1}{2}\pm\frac{2\mu^{2}{\rm H}^{2}+\lambda_{x}}{2\sqrt{4\mu^{2}{\rm H}^{2}(v_{\theta}^{2}q^{2}+\lambda_{x}-2\lambda_{z})+\lambda_{x}^{2}}}}\,$,
in agreement with Eqs. (27) and (28). To determine the relative phase between $\mathcal{Z}^{1}_{q,\alpha}$ and $\mathcal{Z}^{2}_{q,\alpha}$ we write them as
@@
By taking $a=1$, $b=2$ in Eq. (C), we obtain the following equation for the relative phase, $\Delta\varphi_{\alpha}\equiv\varphi_{\alpha}^{1}-\varphi_{\alpha}^{2}$,
$\begin{split}&\sum_{\alpha}\frac{i|\mathcal{Z}_{q,\alpha}^{1}||\mathcal{Z}_{q,\alpha}^{2}|}{\omega_{q,\alpha}\left(\omega^{2}-\omega_{q,\alpha}^{2}\right)}\left[i\,\omega\sin\Delta\varphi_{\alpha}+\omega_{q,\alpha}\cos\Delta\varphi_{\alpha}\right]\\ &\hskip 85.35826pt=-\frac{2\mu{\rm H}\omega}{\det\left[\mathcal{M}(\omega,\bm{q})\right]}\,.\end{split}$ (48)